OpenStudy (anonymous):

Point A is at (3, 2.6) and point B is at (3, 1.4) on a coordinate grid. The distance between the two points is ____. Input numbers and decimal point only, such as 5.8. @DagonsGirl

OpenStudy (anonymous):

@DragonsGirl

OpenStudy (anonymous):

is there a graph ?

OpenStudy (anonymous):

From the Pythagorean Theorem, you get the distance formula between any two points:\[d = \sqrt{\left( x_2-x_1 \right)^2 + \left( y_2-y_1 \right)^2}\]Substitute in the corresponding coordinate for points 1 and 2 and calculate to get the answer.

OpenStudy (anonymous):

ummmm i not really sure this one is hard i would think it would be 2.1 or around 2 units cuz that was i got when i draw it out